# Boundary Attention Constrained Zero-Shot Layout-To-Image Generation

### v1

@@ 4.1 Experimental Setup
Datasets and base models. Following the strategy in the prior studies [27, 16], we perform experiments on a subset of two widely used benchmarks HRS [1] and Drawbench [23]. Specifically, we use four tracks from the HRS benchmark – spatial relationship, size, color, and object counting – which include $1002$, $501$, $501$, and $3000$ text prompts, respectively. These prompts, along with the corresponding layout instructions generated by Chat-4 in [16], are used to evaluate the performance of our method and baseline approaches both quantitatively and qualitatively. Similarly, $39$ text prompts from Drawbench involving spatial and counting specifications are also used in the evaluation. We conduct experiments with the widely-used Stable Diffusion (SD) 1.5 [20] as the base model and perform an ablation study with SD-XL [17] and the fine-tuned L2I model, Gligen [11]. The sampling time step is set to $50$, with the classifier-free guidance weight set to $7.5$. Optimization is applied only to the latent feature $z_{t}$ during the first 10 sampling steps, and $\eta$ is set to 70. Unless specified otherwise, the hyper-parameters $\lambda$ and $\alpha$ are both set to $0.5$.
Baselines and metrics. We compare our proposed scheme, BACON, to six state-of-the-art methods: A&E [2], BoxDiff [28], Layout-Guidance [4], A&R [16], R&B [27], and LoCo [31]. To quantitatively assess our method against these baselines, we use the state-of-the-art object detector, Ground-DINO [13], to detect objects in the synthetic images and predict bounding boxes. These predicted boxes are then compared with the ground truth. We compute precision, recall, and F1 metrics to comprehensively evaluate BACON’s performance in object counting. By comparing the area and centroid of the predicted and ground-truth boxes, we compute the accuracy of object size and spatial relationships. Additionally, Ground-DINO predicts the color of detected objects, which we use to compute color accuracy. The details of the metric computations can be found in the Appendix.
@@ 4.2 Visual Comparisons
In Figure 3, we present visual comparisons between BACON and several competing L2I schemes – Layout, A&R, and R&B – deployed on SD-XL [17]. To thoroughly evaluate the performance of these schemes, we use several complex inputs featuring multiple objects distributed in random positions, which makes it more challenging to generate images with precise object counts. As shown in the figure, BACON demonstrates significant improvement over the baselines in terms of (1) better alignment within the bounding boxes, and (2) more precise object counting. Although the prior works typically generate objects within the corresponding bounding boxes, most of these objects are only partially inside the boxes or even appear in the bounding boxes of other objects due to coarse-grained cross-attention maps, as we discussed earlier. For example, Layout [32] and A&R [16] fail to generate penguins (1st row) within the red bounding boxes and an orange (3rd row) within the blue bounding box. Additionally, none of these prior schemes can generate the correct number of objects according to the prompts and bounding boxes specifications.
To better understand the miscounting failures, we visualize the cross-attention maps obtained from these schemes as shown in Figure 4. Since the cross-attention maps for objects described by the same phrase in the prompt are not independent, existing zero-shot L2I schemes, which aim to encourage cross-attention within the corresponding boxes, often produce overlapping cross-attention maps. For instance, in the RnB scheme, the cross-attention scores are concentrated within three boxes, representing a single eagle, even though three eagles are meant to be depicted. However, boundary-attention loss proposed in BACON can isolate these cross-attention scores to ensure they depict independent objects in the generated images.
@@
Table 1: Quantitative comparisons with baseline schemes. The precision (%), recall (%) and F1 score (%) with respect to object counting are reported in the table. The reported average inference time for generating one image is for a single NVIDIA A100 GPU.
@@
Benchmarks Metrics SD A&E BoxDiff LoCo Layout A&R R&B BACON
@@ 4.3 Quantitative Results
We validate the improvement in spatial controllability and object counting of the proposed BACON framework as compared to existing state-of-the-art zero-shot L2I approaches through quantitative experiments. As shown in Tables 1 and 2, BACON outperforms state-of-the-art baselines on both DrawBench and HRS benchmarks. Notably, BACON demonstrates significant improvements in object counting across three metrics: precision, recall, and F1 score. On DrawBench, BACON outperforms the best baseline, R&B, by $6.34\%$, $7.36\%$, and $6.88\%$ in precision, recall, and F1 score, respectively. On the larger HRS benchmark, BACON consistently achieves the best performance among all L2I schemes, with improvements of $3.78\%$, $0.39\%$, and $1.71\%$ in precision, recall, and F1 score, respectively. Furthermore, BACON surpasses R&B in the spatial relationship metric on both DrawBench and HRS benchmark, with improvement of $5\%$ and $2.59\%$. Additionally, BACON achieves the best performance in color and size metrics, though its advantage over R&B in these areas is relatively small.
Nevertheless, R&B requires to apply Sobel operator [9] to detect edge of cross-attention map for creating minimum bounding rectangle (MBR) which is then used to compute region-aware loss. Therefore, R&B needs more computation in the process of optimizing latent feature $z_{t}$ which leads to longer inference time. As shown in Table. 1, R&B needs more than double time in average for generating one image compared to BACON even though R&B has comparable performance to BACON.
Table 3: Zero-shot schemes can be integrated into existing fully-supervised layout-to-image schemes, e.g., GLIGEN [11]. We compare the improvement of GLIGEN augmented with BACON vs. the competing schemes.
@@
+ BACON 93.22 65.00 83.34 55.76 43.31 42.48
Table 4: The results of experiments on BACON under three settings: (1) $L_{\text{bacon}}$ only includes the region-aware loss (setting R); (2) $L_{\text{bacon}}$ includes both the region-aware loss and boundary-aware loss (setting R + B); (3) using the complete loss proposed in Eq. 12 (setting R + B + Reg).
@@
+ R + B 82.63 53.08 38.12 40.44
+ R + B + Reg 83.34 55.76 43.31 42.28
4.4 Plug and Play with BACON
Although we mainly conduct experiments based on SD 1.5, BACON as well as these prior zero-shot L2I schemes do not assume specific model architecture and can be plugged into arbitrary pre-trained Text-to-Image models using cross-attention blocks. We further evaluate performance of BACON by deploying it on another L2I model, GLIGEN [11], trained by supervised learning with labeled datasets. As shown in Table 3, all zero-shot L2I schemes demonstrate improved performance across nearly all metrics quantifying controllability. Compared to SD 1.5, GLIGEN has the ability to perceive input bounding boxes and incorporate layout information into the prior conditions for sampling the initial noise $z_{0}$. With the initial latent feature $z_{0}$ generated by GLIGEN, the reverse sampler can synthesize an image with a layout that more closely aligns with the given instructions than when using $z_{0}$ produced by SD 1.5. According to Table 3, BACON consistently delivers the best performance in object counting, spatial relationships, and color. While R&B outperforms BACON in the size metric, BACON achieves the second-best performance. In experiments with SD 1.5, plain SD 1.5 generates initial noise $z_{0}$ corresponding to undesirable cross-attention maps that are minimally modified through optimization, resulting in limited improvement of BACON on the size metric. In contrast, GLIGEN generates improved initial noise $z_{0}$ that facilitates BACON’s enhancement.
@@ 4.5 Ablation Study
To investigate the effect of $L_{\text{r}}$, $L_{\text{b}}$ and $L_{\text{reg}}$, the components of BACON’s loss / objective function, we conduct additional experiments using SD 1.5, SD-XL, and GLIGEN under three settings: (1) only the region-attention loss $L_{\text{r}}$ is used (setting R) (2) $L_{\text{r}}$ is combined with the boundary-attention loss $L_{\text{b}}$ (setting R + B); (3) using the complete loss objective as described in Eq. 12 (setting R + B + Reg).
When using only the region-attention loss, BACON synthesizes images similar to those produced by R&B, with the cross-attention map exhibiting the same overlapping issue discussed earlier – see the illustration in Figure 5(a). Adding the boundary-attention loss to the objective resolves the overlapping problem but may cause attention vanishing, as shown in Figure 5(b). This issue does not occur when there is only one bounding box per phrase, due to the constraint imposed by $L_{\text{r}}$. Moreover, the regularization loss $L_{\text{reg}}$ addresses the attention vanishing problem, ensuring that each bounding box corresponds to exactly one object in the generated images, as shown in Figure 5(c).
The quantitative results in Table 4, obtained in experiments on three distinct pretrained diffusion models, further demonstrate the improvement achieved by including $L_{\text{b}}$ and $L_{\text{reg}}$ in the objective function. However, the attention vanishing problem typically arises in settings where multiple objects are assigned adjacent bounding boxes. The text prompts in HRS are relatively simple, so there is minimal difference between R + B and R + B + Reg settings.
Figure 5: Visualization of cross-attention maps of targeting objects generated in three settings: (1) R; (2) R + B; (3) R + B +Reg.
@@ 5 Conclusion
We proposed BACON, a novel zero-shot layout-to-image (L2I) generation scheme utilizing pre-trained diffusion models. We identify two inherent issues that lead to semantic errors and incorrect object counts in generation: (1) coarse-grained cross-attention maps, and (2) overlapping cross-attention. To address these issues, we leverage self-attention maps to refine cross-attention maps and design an objective that combines three distinct loss functions. Comprehensive quantitative and visual results demonstrate that BACON outperforms state-of-the-art L2I schemes and consistently adapts well across varying model architectures.

### v2

@@ 4.1 Experimental Setup
Datasets and base models. Following the strategy in the prior studies [29, 18], we perform experiments on a subset of two widely used benchmarks HRS [2] and Drawbench [24]. Specifically, we use four tracks from the HRS benchmark – spatial relationship, size, color, and object counting – which include $1002$, $501$, $501$, and $3000$ text prompts, respectively. These prompts, along with the corresponding layout instructions generated by ChatGPT-4 in [18], are used to evaluate the performance of our method and baseline approaches both quantitatively and qualitatively. Similarly, $39$ text prompts from Drawbench involving spatial and counting specifications are also used in the evaluation. We conduct experiments with the widely-used Stable Diffusion (SD) 1.5 [22] as the base model and perform an ablation study with SD-XL [19] and the fine-tuned L2I model, Gligen [13]. The sampling time step is set to $50$, with the classifier-free guidance weight set to $7.5$. Optimization is applied only to the latent feature $z_{t}$ during the first 10 sampling steps, and $\eta$ is set to 70. Unless specified otherwise, the hyperparameters $\lambda$ and $\alpha$ are both set to $0.5$.
Baselines and metrics. We compare our proposed scheme, MAC, to six state-of-the-art methods: A&E [3], BoxDiff [30], Layout-Guidance [5], A&R [18], R&B [29], and LoCo [33]. To quantitatively assess our method against these baselines, we use the state-of-the-art object detector, Ground-DINO [15], to detect objects in the synthetic images and predict bounding boxes. These predicted boxes are then compared with the ground truth. We compute precision, recall, and F1 metrics to comprehensively evaluate MAC’s performance in object counting. By comparing the area and centroid of the predicted and ground-truth boxes, we compute the accuracy of object size and spatial relationships. Additionally, Ground-DINO predicts the color of detected objects, which we use to compute color accuracy. The details of the metric computations can be found in the Appendix B.
Figure 3: Visual comparison with different training-free L2I schemes based on Stable Diffusion XL [19]. The layout input and text prompt are shown in the first and second columns, respectively. The layout instructions are also annotated on the generated images with dashed boxes. Our method, MAC, significantly outperforms prior schemes in terms of spatial control and counting accuracy.
@@ 4.2 Visual Comparisons
In Figure 3, we present visual comparisons between MAC and several competing L2I schemes – Layout, A&R, and R&B – deployed on SD-XL [19]. To thoroughly evaluate the performance of these schemes, we use several complex inputs featuring multiple objects distributed in random positions, which makes it more challenging to generate images with precise object counts. As shown in the figure, MAC demonstrates significant improvement over the baselines in terms of (1) better alignment within the bounding boxes, and (2) more precise object counting. Although the prior works typically generate objects within the corresponding bounding boxes, most of these objects are only partially inside the boxes or even appear in the bounding boxes of other objects due to coarse-grained cross-attention maps, as we discussed earlier. For example, Layout [34] and A&R [18] fail to generate penguins (1st row) within the red bounding boxes and an orange (3rd row) within the blue bounding box. Additionally, none of these prior schemes can generate the correct number of objects according to the prompts and bounding boxes specifications.
To better understand the miscounting failures, we visualize the cross-attention maps obtained from these schemes as shown in Figure 4. Since the cross-attention maps for objects described by the same phrase in the prompt are not independent, existing training-free L2I schemes, which aim to encourage cross-attention within the corresponding boxes, often produce overlapping cross-attention maps. For instance, in the R&B scheme, the cross-attention scores are concentrated within three boxes, representing a single eagle, even though three eagles are meant to be depicted. However, marginal-attention loss proposed in MAC can isolate these cross-attention scores to ensure they depict independent objects in the generated images.
@@
Table 1: Quantitative comparisons with baseline schemes. The precision (%), recall (%) and F1 score (%) with respect to object counting are reported in the table.
@@
Benchmarks Metrics SD A&E BoxDiff LoCo Layout A&R R&B MAC
@@ 4.3 Quantitative Results
We validate the improvement in spatial controllability and object counting of the proposed MAC framework as compared to existing state-of-the-art training-free L2I approaches through quantitative experiments. As shown in Tables 2 and 1, MAC outperforms state-of-the-art baselines on both DrawBench and HRS benchmarks. Notably, MAC demonstrates significant improvements in object counting across three metrics: precision, recall, and F1 score. On DrawBench, MAC outperforms the best baseline, R&B, by $6.34\%$, $7.36\%$, and $6.88\%$ in precision, recall, and F1 score, respectively. On the larger HRS benchmark, MAC consistently achieves the best performance among all L2I schemes, with improvements of $3.78\%$, $0.39\%$, and $1.71\%$ in precision, recall, and F1 score, respectively. Furthermore, MAC surpasses R&B in the spatial relationship metric on both DrawBench and HRS benchmark, with improvement of $5\%$ and $2.59\%$. Additionally, MAC achieves the best performance in color and size metrics, though its advantage over R&B in these areas is relatively small.
Nevertheless, R&B requires applying Sobel operator [11] to detect the edge of cross-attention map for creating a minimum bounding rectangle (MBR), which is then used to compute region-aware loss. Therefore, R&B needs more computation in the process of optimizing latent feature $z_{t}$, which leads to longer inference time. As shown in Table. 2, R&B needs more than double time on average for generating one image compared to MAC, even though R&B has comparable performance to MAC.
Table 3: Zero-shot schemes can be integrated into existing fully-supervised layout-to-image schemes, e.g., GLIGEN [13]. We compare the improvement of GLIGEN augmented with MAC vs. the competing schemes.
@@
+ MAC 93.22 65.00 83.34 55.76 43.31 42.48
Table 4: The results of experiments on MAC under three settings: (1) $L_{\text{mac}}$ only includes the region-aware loss (setting R); (2) $L_{\text{mac}}$ includes both the region-aware loss and marginal-aware loss (setting R + M); (3) using the complete loss proposed in Eq. 12 (setting R + M + Reg).
@@
+ R + M 82.63 53.08 38.12 40.44
+ R + M + Reg 83.34 55.76 43.31 42.28
4.4 Plug and Play with MAC
Although we mainly conduct experiments based on SD 1.5, MAC, as well as these prior training-free L2I schemes do not assume specific model architecture and can be plugged into arbitrary pre-trained Text-to-Image models using cross-attention blocks. We further evaluate the performance of MAC by deploying it on another L2I model, GLIGEN [13], trained by supervised learning with labeled datasets. As shown in Table 3, all training-free L2I schemes demonstrate improved performance across nearly all metrics quantifying controllability. Compared to SD 1.5, GLIGEN has the ability to perceive input bounding boxes and incorporate layout information into the prior conditions for sampling the initial noise $z_{0}$. With the initial latent feature $z_{0}$ generated by GLIGEN, the reverse sampler can synthesize an image with a layout that more closely aligns with the given instructions than when using $z_{0}$ produced by SD 1.5. According to Table 3, MAC consistently delivers the best performance in object counting, spatial relationships, and color. While R&B outperforms MAC in the size metric, MAC achieves the second-best performance. In experiments with SD 1.5, plain SD 1.5 generates initial noise $z_{0}$ corresponding to undesirable cross-attention maps that are minimally modified through optimization, resulting in limited improvement of MAC on the size metric. In contrast, GLIGEN generates improved initial noise $z_{0}$ that facilitates MAC’s enhancement.
@@ 4.5 Ablation Study
To investigate the effect of $L_{\text{r}}$, $L_{\text{b}}$ and $L_{\text{reg}}$, the components of MAC’s loss / objective function, we conduct additional experiments using SD 1.5, SD-XL, and GLIGEN under three settings: (1) only the region-attention loss $L_{\text{r}}$ is used (setting R); (2) $L_{\text{r}}$ is combined with the marginal-attention loss $L_{\text{m}}$ (setting R + M); (3) using the complete loss objective as described in Eq. 12 (setting R + M + Reg).
When using only the region-attention loss, MAC synthesizes images similar to those produced by R&B, with the cross-attention map exhibiting the same overlapping issue discussed earlier – see the illustration in Figure 5(a). Adding the marginal-attention loss to the objective resolves the overlapping problem but may cause attention vanishing, as shown in Figure 5(b). This issue does not occur when there is only one bounding box per phrase, due to the constraint imposed by $L_{\text{r}}$. Moreover, the regularization loss $L_{\text{reg}}$ addresses the attention vanishing problem, ensuring that each bounding box contains exactly one object in the generated images, as shown in Figure 5(c).
The quantitative results in Table 4, obtained in experiments on three distinct pretrained diffusion models, further demonstrate the improvement achieved by including $L_{\text{m}}$ and $L_{\text{reg}}$ in the objective function. However, the attention vanishing problem typically arises in settings where multiple objects are assigned adjacent bounding boxes. The text prompts in HRS are relatively simple, so there is minimal difference between R + M and R + M + Reg settings.
Figure 5: Visualization of cross-attention maps of targeting objects generated in three settings: (1) R; (2) R + M; (3) R + M +Reg.
@@ 5 Conclusion
We proposed MAC, a novel training-free layout-to-image (L2I) generation scheme utilizing pre-trained diffusion models. We identify two inherent issues that lead to semantic errors and incorrect object counts in generation: (1) coarse-grained cross-attention maps, and (2) overlapping cross-attention. To address these issues, we leverage self-attention maps to refine cross-attention maps and design an objective that combines three distinct loss functions. Comprehensive quantitative and visual results demonstrate that MAC outperforms state-of-the-art L2I schemes and consistently adapts well across varying model architectures. We envision our MAC as a milestone that can inspire the follow-up research in training-free layout-to-image generation.
Appendix A: Experimental Details
In this section, we report the model architectures and hyper-parameters used in the experiments across all layout-to-image schemes. We employed the DDIM scheduler [26] with $50$ denoising steps in all experiments, performing latent variable optimization only within the first 10 steps, with a maximum of 5 iterations per step. The step size for updating the latent variable $z_{t}$ was set to $70$, and the loss threshold value for early stopping was set to $10^{-6}$. In all experiments, the weight for classifier-free guidance was set to $7.5$. In the experiments with MAC, $\lambda$ and $\alpha$ for combining boundary-attention loss and regularization loss were both set to $0.25$.
Appendix B: Evaluation Details
We employed the state-of-the-art GroundDINO [15] to detect objects in the synthetic images generated with the input prompts and layout instruction from DrawBench and HRS benchmarks. Specifically, GroundDINO generates multiple predicted bounding boxes corresponding to predicted categories on the synthetic images with threshold value for confidence $0.25$. Those boxes are then used to compute various metrics for measuring the spatial controllability and counting accuracy.
Object Counting
To evaluate the layout-to-image schemes in object counting, we record the number of predicted bounding boxes $n^{(i)}_{\text{pred}}$ corresponding to the phrase $\mathbf{p}^{(i)}$, and compute the correct number of boxes and false number of boxes by
$n^{(i)}_{\text{cor}}=\min(n_{\text{pred}}^{(i)},n^{(i)}_{\text{gt}})$, (13)
$n^{(i)}_{\text{fal}}=\max(n_{\text{pred}}^{(i)}-n^{(i)}_{\text{gt}},0),$ (14)
$n^{(i)}_{\text{neg}}=\max(n^{(i)}_{\text{gt}}-n_{\text{pred}}^{(i)},0),$ (15)
where $n^{(i)}_{\text{gt}}$ is the ground-truth number. With $n^{(i)}_{\text{cor}}$ and $n^{(i)}_{\text{fal}}$, we can obtain
$\emph{precision}=\frac{\sum_{i\in\mathcal{I}}n^{(i)}_{\text{cor}}}{\sum_{i\in\mathcal{I}}n^{(i)}_{\text{cor}}+\sum_{i\in\mathcal{I}}n^{(i)}_{\text{fal}}}$, (16)
$\emph{recall}=\frac{\sum_{i\in\mathcal{I}}n^{(i)}_{\text{cor}}}{\sum_{i\in\mathcal{I}}n^{(i)}_{\text{cor}}+\sum_{i\in\mathcal{I}}n^{(i)}_{\text{neg}}}$, (17)
where $\mathcal{I}$ denotes indices of the phrases in the prompt, which is defined in the Section 3.1 in the main paper. With the metrics of precision and recall, we compute F1 score by
$\emph{F1}=\frac{2*\emph{precision}*\emph{recall}}{\emph{precision}+\emph{recall}}.$ (18)
Spatial Accuracy
In the experiments on spatial relationship, phrases in the input prompts are given grouth-truth relationship. For instance, in the prompt a cat is on the left of a dog, the two phrases cat and dog have the spatial relationship on the left of. By comparing the mean points of the predicted boxes for cat and dog as
$(\mathbf{x}_{1},\mathbf{y}_{1})=(\frac{\mathbf{x}_{1}^{\text{left}}+\mathbf{x}_{1}^{\text{right}}}{2},\frac{\mathbf{y}_{1}^{\text{left}}+\mathbf{y}_{1}^{\text{right}}}{2}),$ (19)
$(\mathbf{x}_{2},\mathbf{y}_{2})=(\frac{\mathbf{x}_{2}^{\text{left}}+\mathbf{x}_{2}^{\text{right}}}{2},\frac{\mathbf{y}_{2}^{\text{left}}+\mathbf{y}_{2}^{\text{right}}}{2}),$ (20)
and record the number of correct prediction
[…]
and compute the spatial accuracy
$\emph{ACC}_{\text{spatial}}=\frac{\sum_{i\in\mathcal{I}}n_{\text{cor}}^{(i)}}{|\mathcal{I}|}.$ (22)
Size Accuracy
In the experiments on size relationship, phrases in the input prompts are given grouth-truth relationship. For instance, in the prompt a cat is smaller than a dog, the two phrases cat and dog have the size relationship smaller. By comparing the area of the predicted boxes $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ for cat and dog as
[…]
and compute the size accuracy
$\emph{ACC}_{\text{size}}=\frac{\sum_{i\in\mathcal{I}}n_{\text{cor}}^{(i)}}{|\mathcal{I}|}.$ (24)
Figure 6: Example of overlapping objects in a box.
Color Accuracy
In the experiments on color, phrases in the input prompts are given color instruction. For instance, in the prompt a white cat and a black dog, the two phrases cat and dog have the color white and black. Similarly, we can compute the color accuracy:
[…]
and compute the color accuracy
$\emph{ACC}_{\text{color}}=\frac{\sum_{i\in\mathcal{I}}n_{\text{cor}}^{(i)}}{|\mathcal{I}|}.$ (26)
Complex Layouts
As shown in Figure 6, our method excels at handling complex and crowded scenes compared to previous approaches, thanks to the boundary-attention constraint. For overlapping objects, the user can assign a unique bounding box to each object and use a distinct phrase to describe them. For example: “One penguin is standing on the left of the beach, and five baby penguins are on the right.” The user can assign one box for the phrase One penguin while one box for the phrase baby penguins, as illustrated on the top right. However, the cross-attention maps for multiple objects within the same box inevitably overlap, leading to incorrect counting. We recognize that these training-free L2I schemes, including ours, may struggle with accurate counting in overlapping boxes, resulting in unsatisfactory generation in crowded scenarios. Addressing this limitation would be interesting for future work.
Hyper-parameter Sensitivity
The results in Table 4 highlight the roles of the three loss functions: (1) region-attention loss regulates object location, (2) boundary-attention loss improves object counting accuracy, and (3) regularization loss prevents the attention map from vanishing. Therefore, we set $\lambda=\alpha=0.5<1$ as location regulation is the highest priority. To provide further insights, we will conduct additional experiments with varying $\lambda$ and $\alpha$ in the revised version.
Text-Layout Mismatch
The entire discussion on improving object counting is based on the assumption that the user inputs a consistent layout and text prompts, which is a reasonable assumption, our method will not improve the counting otherwise. Under the same assumption, the existing schemes have shown poor performance in object counting accuracy, as illustrated in Table 1, because it is challenging to accurately determine the count relying solely on the text prompt.